I have 5g of a power which has an Mr of 164.16 and a solubility of 8.2mg/mL in DMSO. i would like to dilute the powder into the DMSO and obtain concentrations of 5mM, 20mM and 50nM. What formula do i need to use and how would i work this out.
1
Expert's answer
2019-11-12T12:01:58-0500

Find moles of powder:

"n=\frac{m}{M} =\frac{5}{164.16} = 0.03 mol"


1) 5mM solution

Find the volume of 5 mM solution when we dissolve all 0.03 moles in DMSO:


"V=\frac{n}{c} = \frac{0.03mol}{0.005 \frac{mol}{l}}= 6.092 l = 6092 ml"

Check up if all 5 g of powder will dissolve in 6 ml of solution:

0.0082 g of powder dissolve in 1 ml

x g of powder dissolve in 6092 g

Solve the proportion:


"\frac{0.0082}{x} =\frac{1}{6092}"

"x= 49.95 g"

So we have less grams of powder then 49.95 , then all 5 grams will dissolve.

To make 5 mM solution we should take 5 g of powder and add enough DMSO to reach the line on the beaker of 6092 ml.


2) 20mM solution


"V=\frac{0.03 mol}{0.020 \frac{mol}{l}} = 1.5 l =1500 ml"

Check up if all 5 g of powder will dissolve in 1500 ml of solution:

0.0082 g of powder dissolve in 1 ml

x g of powder dissolve in 1500 g

Solve the proportion:


"\frac{0.0082}{x} =\frac{1}{1500}"

"x = 12.3 g"

So we have less grams of powder than 12.3 , then all 5 grams will dissolve.

To make 20 mM solution we should take 5 g of powder and add enough DMSO to reach the line on the beaker of 1500 ml.


3) 50nM solution

We can see that the concentration of this solution is 10000 times less than of the first solution "\frac{50\times10^{-9}}{5\times10^{-3}} = 1\times 10^{-5}"

So we need to take 1 ml of the first solution and to dilute it with DMSO 10000 times.